Pigment of chicken feathers is regulated by two genes, the gene for feather pigment C, and a gene that inhibits pigment production, I. A mother chicken, who is homozygous dominant for gene C, and heterozygous for gene I, is crossed with a father chicken, who is heterozygous for the pigment gene, and heterozygous for gene I. Inhibition is dominant to uninhibited, pigmented feathers are dominant to unpigmented feathers. The inhibitor gene (I) will prevent expression of pigmented chicken feathers.

A) What are the potential haplotypes of the gametes produced by the mother for these two genes? Use the letters C, c, I, and i for the genes. C is for color and I is for inhibitor Enter your answer as a list of haplotypes separated by spaces such as AB ab Ab aB. Do not write a sentence. Do not list duplicate haplotypes. Alphabatize your haplotypes, e.g. Ab, not bA.

B) What are the potential haplotypes of the gametes produced by the father for these two gametes? Use the letters C, c, I, and i for the genes. C is for color and I is for inhibitor Enter your answer as a list of haplotypes separated by spaces such as AB ab Ab aB. Do not write a sentence. Do not list duplicate haplotypes. Alphabetize your haplotypes, e.g. Ab, not bA.

C) What is the probability of one of the offspring of these two chickens having pigmented feathers? % Enter your answer as a percent, but do not include the percent symbol, e.g. 6.25, not 6.25%, 0.0625, or 1/16 Report your answer to one decimal place, e.g. 66.7, not 66.66 or 66.60.

D) In a population of 150 chickens, produced from the same parents, how many chickens would you expect to have white feathers? Round your answer to the lowest whole number since you cannot have a partial chicken, e.g. 65.8 would round to 65, as would 65.4, 65.5.
Answer #1

Genotype of mother - CCIi

Genotype of father - CcIi

potential haplotypes of the Gametes produced by mother-

CI Ci

Potential haplotypes of the gametes produced by father-

CI Ci cI ci

CI Ci
CI CCII CCIi
Ci CCIi CCii
cI CcII CcIi
ci CcIi Ccii

Number of offaprings with white feather= 2/8= 0.25

Out of 150, total number of offsprings with white feathers= 2/8 X 150= 37.5 i. e. 37